Convert
12x10^34 Å to Km
48.5x10^5nM to mM
233.6x10^-20 PA to nA
1
Expert's answer
2019-08-19T09:51:18-0400
"12\times 10^{34}\text{\AA}=12\times 10^{24}\:\text{m}=12\times 10^{21}\:\text{km}"

"48.5\times 10^{5}\:\text{nm}=48.5\times 10^{-4}\:\text{m}=4.85\:\text{mm}"

"233.6\times 10^{-20}\:\text{pA}=233.6\times 10^{-32}\:\text{A}=233.6\times 10^{-23}\:\text{nA}"